# HAMMER: Heterogeneous, Multi-Robot Semantic Gaussian Splatting

### v1

@@ Abstract
3D Gaussian Splatting offers expressive scene reconstruction, modeling a broad range of visual, geometric, and semantic information. However, efficient real-time map reconstruction with data streamed from multiple robots and devices remains a challenge. To that end, we propose HAMMER, a server-based collaborative Gaussian Splatting method that leverages widely available ROS communication infrastructure to generate 3D, metric-semantic maps from asynchronous robot data-streams with no prior knowledge of initial robot positions and varying on-device pose estimators. HAMMER consists of (i) a frame alignment module that transforms local SLAM poses and image data into a global frame and requires no prior relative pose knowledge, and (ii) an online module for training semantic 3DGS maps from streaming data. HAMMER handles mixed perception modes, adjusts automatically for variations in image pre-processing among different devices, and distills CLIP semantic codes into the 3D scene for open-vocabulary language queries. In our real-world experiments, HAMMER creates higher-fidelity maps (2x) compared to competing baselines and is useful for downstream tasks, such as semantic goal-conditioned navigation (e.g., “go to the couch”). Accompanying content available at hammer-project.github.io.
@@
Figure 1: HAMMER takes streaming data from a heterogeneous team of robot-mounted sensors (e.g. ZED camera) and edge devices (e.g. Aria glasses [1]). The robots and devices run on-board feature-based SLAM in unaligned local frames. On the server side, HAMMER aligns data streams into a global frame. Aligned data streams are integrated into an online, collaboratively constructed $3$DGS map. The map can be leveraged for open-vocabulary object localization, and planning.
@@ I Introduction
In this work, we propose HAMMER, Heterogeneous Asynchronous Multi-robot Mapping of Environmental Radiance. HAMMER enables a team of robots to collaboratively construct a 3DGS map of an unknown environment. Under HAMMER, robots broadcast color images and geometric sensing data, posed using the robot’s on-board localization system, to a mapping server. Semantic information is extracted from the images and the local poses of each robot are aligned to a global map frame. A semantic $3$DGS map is trained online using the processed data streams from all robots. A server-based architecture incurs negligible requirements on on-board computational resources and can be deployed wherever typical communication infrastructure (e.g. WiFi/5G) is available. Importantly, HAMMER does not require any prior knowledge on the initial robot positions or inter-robot communication/observations, enabling asynchronous and ad-hoc robot deployments.
HAMMER is designed to generalize to a wide range of robots and devices, combining the advantages of each device into a single map. A shared map enables these robots to have more comprehensive spatial awareness compared to their own local maps. For example, drones can cover blind spots of mobile ground robots performing semantically-guided navigation. We intentionally propose a device-server architecture with a centralized 3DGS training module to ensure that HAMMER is real-time executable with existing robots and wearable devices. Computing requirements for 3DGS training are currently beyond the on-board compute capabilities of most robots and wearables. When on-board computing matches the power of current high-end GPUs, fully distributed algorithms [11] may become a compelling alternative.
We validate the performance of HAMMER on real-world streaming data from a heterogeneous team of egocentric human-wearable devices (Aria glasses [1]) and ground robots with a stereo camera. Furthermore, we demonstrate the practical effectiveness of HAMMER by simultaneously optimizing the $3$DGS map and performing multi-robot trajectory optimization with open vocabulary, semantic goal specification. The contributions of HAMMER (Figure 1) are:
  • a robust and metric inter-robot frame alignment using RGB images to align coordinate systems across devices with different image sensors and SLAM algorithms,
  • and an online training scheme for $3$DGS from heterogeneous, asynchronous devices.
We demonstrate HAMMER in hardware deployments with 3-4 devices across 2 real-world environments. HAMMER provides more than 40% better Mean-Squared-Error (MSE) compared to the Di-NeRF [12] baseline.
@@ II Related Work
Radiance Fields and Gaussian Splatting. NeRFs, proposed in [2], use neural networks to model radiance fields from posed RGB images and demonstrate photorealistic novel view synthesis and high-fidelity geometric reconstruction by optimizing these networks using a photometric loss with a differentiable rendering process. Follow up works showed that more sophisticated ray-sampling methods [13] greatly improves reconstruction accuracy, and optimized model architectures [14] can reduce training times by orders of magnitude. However, NeRFs remain computationally expensive to optimize and query. As an implicit representation, NeRFs make it difficult to extract the underlying scene geometry. On the other hand, $3$DGS [3] maintains many of the benefits of NeRFs and is significantly faster to optimize. The use of geometric primitives as a representation also facilitates collision quantification and contact modeling with the environment [7].
@@
Online 3DGS Training and 3DGS-based SLAM. $3$DGS is typically optimized using the following offline optimization scheme: (i) image capture using a high-resolution camera, (ii) image pose estimation using structure-from-motion (SfM) like COLMAP [15], and (iii) $3$DGS map optimization on the static dataset. However, this processing and training scheme is not well suited for robotics tasks where data is incrementally received from onboard sensors and prevents the $3$DGS map from being used during deployment. LEGS [10] proposes using an incrementally optimized semantic 3DGS map with camera poses initialized from an on-board SLAM system. Gaussian Splatting SLAM [4] and SplaTAM [16] go a step further, using $3$DGS as the map representation for SLAM. While LEGS does discuss using multiple cameras, these cameras are on a fixed rigid frame, and does not constitute a multi-robot system.
@@
Multi-robot Mapping. Mapping using multi-robot systems is a widely studied field, and many diverse methods have been proposed. Historically, multi-robot mapping and SLAM systems have utilized more traditional robot map representations like point clouds [17] and meshes [18]. Collaborative mapping is well-suited for scenarios requiring robust and rapid exploration [19].
MACIM [20] proposes a decentralized multi-robot mapping algorithm that optimizes a signed distance function, while other approaches use NeRFs [21, 12]. Of these works, only Di-NeRF [12] addresses the problem of frame alignment from the robots’ local frames to a global map frame, which it does through gradient-based optimization. MULAN-WC [22] proposes using inter-robot observations from wireless signal measurements to align robots for multi-robot NeRF training. In contrast, our work does online continual inter-robot frame alignment from multiple heterogeneous devices.
Developed concurrently to this work, MAGiC-SLAM [23] proposes a multi-robot $3$DGS SLAM pipeline that enables robots that are running independent $3$DGS SLAM instances to fuse their individual maps into a global map. The key differences when compared to HAMMER are that MAGiC-SLAM relies on the capability of individual devices to train local $3$DGS maps, and that it does not account for heterogeneous sensing devices. There is no publicly available code for this method.
III Problem Setup
@@
We consider a set of $N$ robots, indexed $i\in\{1,\dots,N\}$, deployed in a shared environment. Their deployments can be asynchronous, potentially lacking any temporal overlap between individual robot deployments. Each robot produces color images, geometric information (e.g. depth images or point clouds), and camera pose estimates in $SE(3)$ with respect to an arbitrary local coordinate frame $\mathcal{T}^{i}$. These estimates can be retrieved from on-board VIO or SLAM commonly available on commercial robots and devices. The robots continuously stream all data to a mapping server during deployment. The data is received incrementally by the server as the robot navigates through the environment. The goal is to continuously incorporate every robot’s data into the $3$DGS map within a consistent global coordinate frame $\mathcal{T}^{g}$. We assume there is sufficient visual overlap within the data streams of the robots such that alignment is possible.
Figure 2: An illustration of HAMMER’s inter-robot alignment module. Images from aligned and unaligned robots are matched using a place recognition feature extractor. When a correspondence is verified, images from the two robots are grouped together and passed to a localized SfM pipeline that estimates the relative pose between the images. The SfM poses are then used to find a single transform to bring the unaligned robot’s local frame into the aligned robot’s global frame. Images and transformed SLAM poses from all robots are then streamed into the data loader for $3$DGS training.
IV Method
We decompose the problem into (i) inter-robot alignment where the data from all robots is aligned to a common reference frame, and (ii) multi-robot $3$DGS mapping where streaming data is incorporated into a $3$DGS map.
IV-A Inter-robot Alignment
In HAMMER’s inter-robot alignment module (see Figure 1) we begin by selecting one of the contributing robots as the origin, and making its local reference frame the global frame, $\mathcal{T}^{g}$. This robot may be selected at random or simply as the first robot that connects to the server. The selected reference frame is assumed to be scaled to metric units. This origin robot immediately begins integrating its streamed data into the 3DGS map. Already connected, or newly connected, robots that are not the origin robot are designated as unaligned and stream their data to a cache that will be used to align the robots to the global frame and then aggregated into the 3DGS optimization data pool (see Section IV-B). The alignment module runs at fixed intervals—parallel to the mapping process—as new data is received from the unaligned robots.
IV-A1 Correspondences
@@
In the first stage of the alignment process, HAMMER detects potential candidate correspondences between the data of the unaligned robots and that of the aligned robots. Regardless of alignment status, a place-recognition feature vector is extracted for every color image received from the collaborating robots using NETVLAD [24]. Additionally, the features from the unaligned robots are periodically compared to those of the aligned robots using cosine similarity. The similarities are thresholded with $\gamma_{ij}$ where $i,j$ indicates the index of the aligned and unaligned robot, respectively. For pairs of images with similarities exceeding $\gamma_{ij}$, the alignment module performs an additional feature-matching verification step where SuperPoint local features [25] are extracted for each image, and then matched using LightGlue [26]. If the fraction of features matched during verification exceeds a fixed threshold (0.25 in our experiments) then the image pair is accepted as a potential inter-robot correspondence.
IV-A2 Adaptive Thresholding
@@
When no image pairs pass the threshold, then $\gamma_{ij}$ is increased to match the mean of all place recognition similarities greater than the previous value. This thresholding scheme is used because multiple instances of the same camera (e.g. two robots with ZED cameras) tend to have higher similarity regardless of visual content, resulting in spurious matches. Dynamic thresholding allows initial values of $\gamma_{ij}$ to be set to a low value for all robot pairs, circumventing any hyper-parameter tuning that would otherwise be necessary. While this procedure for determining candidate alignment image pairs grows quadratically in the number of images per robot, many of these operations can be cached, forming a relatively small computational footprint.
IV-A3 Localized Structure-from-Motion
@@
For every potential correspondence between an aligned and unaligned robot, a small set of color images within some fixed temporal window size $W=16$ is selected from each robot’s data cache on the server. Subsets of the pair’s data streams are used to perform localized SfM, as opposed to bulk SfM where the whole data stream is used. We use the COLMAP backend [15] with SuperPoint features and the SuperGlue matcher [27], which demonstrates robustness in aligning images from different device types. Importantly, the localized SfM does not use any pose information from the robot’s SLAM systems, producing an independent relative pose estimate between robots.
IV-A4 Registration
The output of the localized SfM procedure lies in an arbitrary coordinate frame, $\mathcal{T}^{s}$. The final stage of alignment determines the scaled transformation mapping poses from the local frame of the unaligned robot, $\mathcal{T}^{j}$, to the global frame of the aligned robot, $\mathcal{T}^{g}$. The relative scaled transform, rotation $R^{jg}\in SO(3)$, translation $t^{jg}\in\mathbb{R}^{3}$, and scaling $s^{jg}\in\mathbb{R}^{+}$, transforms a pose $(R^{j},t^{j})$ in $\mathcal{T}^{j}$ to $\mathcal{T}^{g}$ through
$t^{g}$ $=s^{jg}R^{jg}t^{j}+t^{jg},\;\;R^{g}=R^{jg}R^{j},$ (1)
where $(R^{g},t^{g})$ is the transformed pose. Scaled transformations account for non-metric localization software used by any robot and the SfM intermediary frame. So long as the origin frame is set using a robot with a metric localization system, the map will retain the metric scale.
@@ IV-A4 Registration
Computing the relative transformation for unaligned robots requires solving two instances of a modified absolute orientation problem [28]. One instance computes the scaled transformation from the unaligned $\mathcal{T}^{j}$ to the intermediary SfM frame $\mathcal{T}^{s}$. The second instance transforms poses from $\mathcal{T}^{s}$ to the global frame $\mathcal{T}^{g}$. The first instance ingests an ordered sequence of poses $\{(R_{k}^{j},t_{k}^{j})\}_{k=1}^{W}$ from the unaligned local frame and $\{(R_{k}^{s},t_{k}^{s})\}_{k=1}^{W}$ from the corresponding poses in the localized SfM solution.
To compute these transformations we solve the following rotation-aware absolute orientation problem
@@
where $||\cdot||_{F}$ denotes the Frobenius norm. Equation 2 optimizes the scaling, rotation, and translation $(s,R,t)$ between the two frames with a small regularization term on the rotation to break ties. For example, pose sequences where both datasets progress linearly (like for ground robots) may not have a unique rotation that minimizes the first term of (2). Therefore, the second term serves to optimize over the remaining degrees of freedom of $R$ subject to some small $\epsilon=0.01$. While no closed form solution exists for (2), each individual term has a closed form solution dependent on an expression of the form $SVD(H)$. Therefore, we can compute an approximately optimal solution dependent on $SVD(H_{1}+\epsilon H_{2})$.
The result of approximately solving (2) is the scaled transformation from the unaligned $\mathcal{T}^{j}$ frame to the localized SfM $\mathcal{T}^{s}$ frame. Repeating the same procedure between poses in the global frame set by the aligned robots and their corresponding poses in the SfM solution produces the transformation from $\mathcal{T}^{s}$ to $\mathcal{T}^{g}$. Finally, the composition of both transformations yields $(s^{jg},R^{jg},t^{jg})$ the relative frame transformation from the unaligned robot’s local coordinate frame to the global coordinate frame.
The proposed inter-robot alignment module is robust and fails gracefully by terminating early or rejecting spurious alignments. These events can occur during image-correspondence verification, after localized SfM if the solution fails to converge, and during the absolute orientation procedures if the solution residuals are high. Rejection at any of these stages avoids spurious alignments due to on-board localization errors or failures during the localized SfM solve.
@@
Figure 2 illustrates a high-level overview of the entire alignment process. Importantly, the process treats the on-board localization algorithms as black-boxes and only uses the resulting camera poses and color images as input. This attribute of HAMMER dramatically simplifies the integration of new devices due to the ubiquity and variety of on-board localization frameworks. In contrast, alignment methods that require access to lower-level information from each individual robot’s localization system (e.g. pose graphs or sparse maps) will fail if the software is closed-source.
Figure 3: Meshes extracted from the $3$DGS map trained with HAMMER during runtime (left) and the final map reconstruction (right). The attribution mesh coloring is based on portion of the map which was contributed by each robot.
IV-B Multi-robot 3DGS Mapping

### v2

@@ Abstract
3D Gaussian Splatting offers expressive scene reconstruction and can model a broad range of visual, geometric, and semantic information. However, efficient real-time map reconstruction with data streamed from multiple robots and devices remains a challenge. To that end, we propose HAMMER, a server-based multi-robot Gaussian Splatting method that leverages ROS communication infrastructure to generate 3D, metric-semantic maps from asynchronous robot data-streams. HAMMER consists of (i) a one-time frame alignment module that transforms local SLAM poses and image data into a global frame and requires no prior relative pose knowledge, and (ii) an online module for continually training semantic 3DGS maps from streaming data. HAMMER handles mixed perception modes, adjusts automatically for variations in image pre-processing among different devices, and distills CLIP semantic codes into the 3D scene for language queries. In real-world experiments, HAMMER creates better maps compared to baselines and is useful for downstream tasks, such as semantic navigation (e.g., “go to the couch”). Accompanying content at hammer-project.github.io.
Index Terms:
Mapping, Multi-robot Systems, Visual Learning
@@
Figure 1: HAMMER takes streaming image and depth data from a heterogeneous team of robots and edge devices (e.g. Aria Glasses [1]), each running on-board SLAM in unaligned local frames. On a GPU server, HAMMER aligns new device data streams into a global frame in a one-time computation (1). All aligned data streams are then integrated into an online continually trained semantic $3$DGS map (2). The map can be leveraged for open-vocabulary object localization and planning.
@@ I Introduction
In this work, we propose HAMMER, Heterogeneous Asynchronous Multi-robot Mapping of Environmental Radiance. HAMMER enables a server communicating with a team of robots to construct a joint 3DGS map of an unknown environment. First, robots broadcast color images and depth maps, annotated with poses using the robot’s on-board localization system, to a mapping server. The server extracts semantic information from the images, aligns the local poses of each robot to a global map frame, and continually trains in real-time a metric-semantic $3$DGS map using the robot data streams.
A server-based architecture allows our method to be used with existing robot and edge device hardware without high-powered GPUs, while leveraging typical communication infrastructure (e.g. WiFi/5G) to offload demanding frame alignment and 3DGS training to a central GPU server. We intentionally propose this partially centralized architecture to ensure that HAMMER is real-time executable with existing robots, wearable devices, and WiFi/5G communication infrastructure. Computing requirements for 3DGS training are currently beyond the on-board compute capabilities of most robots and wearables. When on-board computing matches the power of current high-end GPUs, fully distributed algorithms [11, 12] may become a compelling alternative.
HAMMER is designed to generalize to a wide range of robots and devices, combining the advantages of each device into a single map. A shared map enables these robots to have comprehensive spatial awareness compared to their own local maps. For example, drones can cover blind spots of mobile ground robots performing semantically-guided navigation.
The contributions of HAMMER (Fig. 1) are (1) a robust and metric inter-robot frame alignment using RGB images to align coordinate systems across devices with different image sensors and SLAM algorithms, and (2) an online $3$DGS training scheme from heterogeneous, asynchronous devices that produce posed RGB images with depth or point clouds. The frame alignment (1) occurs only once when a new robot joins the map server, while the real time training (2) operates continually, integrating aligned robots’ data into the server map as it is streamed. Importantly, HAMMER does not require any prior knowledge on the initial robot poses nor inter-robot observations, enabling asynchronous and ad-hoc deployments.
We compare HAMMER against state-of-the-art multi-robot mapping baselines [13, 14] across four scenes from the Replica dataset[15], where HAMMER outperforms or matches the baseline while requiring less than a tenth the computation time. Furthermore, in hardware deployments with 3-4 devices across two real-world environments, HAMMER provides more than 40% better Mean-Squared-Error for novel-view image reconstruction, compared to [16].
@@ II Related Work
Radiance Fields and Gaussian Splatting. NeRFs, proposed in [2], use neural networks to model radiance fields from posed RGB images and demonstrate photorealistic novel view synthesis and high-fidelity geometric reconstruction by optimizing these networks using a photometric loss with a differentiable rendering process. Follow up work with optimized model architectures [17] greatly reduced NeRF training times. However, NeRFs remain computationally expensive to optimize and query. As an implicit representation, NeRFs make it difficult to extract the underlying scene geometry. On the other hand, $3$DGS [3] maintains many of the benefits of NeRFs and is significantly faster to optimize. The use of geometric primitives as a representation also facilitates collision and contact modeling with the environment [7].
@@
Online 3DGS Training and 3DGS-based SLAM. $3$DGS is typically optimized using the following offline optimization scheme: (i) images are captured using a high-resolution camera, (ii) poses are estimated using structure-from-motion (SfM) like COLMAP [18], and (iii) a $3$DGS map is optimized on the static dataset. However, this processing and training scheme is not well suited for robotics tasks where data is incrementally received from onboard sensors and prevents the $3$DGS map from being used during deployment. LEGS [10] proposes using an incrementally optimized semantic 3DGS map with camera poses initialized from an on-board SLAM system. Both [4] and [19] go a step further, using $3$DGS as the map representation for SLAM.
@@
Multi-robot Mapping. Mapping using multi-robot systems is a widely studied field, and many diverse methods have been proposed. Historically, multi-robot mapping and SLAM systems have utilized more traditional robot map representations like point clouds [20, 14] and meshes [21]. Collaborative mapping is well-suited for scenarios requiring robust and rapid exploration [22]. MACIM [23] proposes a decentralized multi-robot mapping algorithm that optimizes a signed distance function, while other approaches use NeRFs [24, 16]. Of these works, only Di-NeRF [16] addresses the problem of inter-robot frame alignment, which it does through gradient-based optimization.
Developed concurrently to this work, MAGiC-SLAM [13] proposes a multi-robot $3$DGS SLAM pipeline comprised of robots running independent $3$DGS SLAM instances whose individual maps are fused into a global map. MAGiC-SLAM requires significantly greater computational demands compared to HAMMER. Either each robot is equipped with a GPU capable of 3DGS optimization or data is streamed to a server with $N+1$ GPUs, where $N$ is the number of concurrently operating robots. Instead, HAMMER only requires a server with a single GPU and does not impose additional compute constraints on the robots beyond requiring a stream of pose and image/depth data. HAMMER offers similar or better reconstruction quality at one tenth the computation time, is built on ROS for real robots and on-line continual mapping, and embeds semantic codes for language based map queries.
III Method
III-A Problem Setup
@@
We consider a set of $N$ robots, indexed $i\in\{1,\dots,N\}$, deployed in a shared environment. Their deployments can be asynchronous, potentially lacking any temporal overlap between individual robot deployments. Each robot produces color images, geometric information (e.g. depth images or point clouds), and camera pose estimates in $SE(3)$ with respect to an arbitrary local coordinate frame $\mathcal{T}^{i}$. These estimates can be retrieved from onboard VIO or SLAM commonly available on commercial robots and devices. The robots continuously stream data to a mapping server during deployment. The data is received incrementally by the server as the robot navigates through the environment. The goal is to continuously incorporate every robot’s data into the $3$DGS map within a consistent global coordinate frame $\mathcal{T}^{g}$. We assume that there is sufficient visual overlap within the data streams of the robots such that alignment is possible. We also assume that at least one of the robots provides pose estimates in a metric coordinate frame.
III-B Inter-robot Alignment
When a new robot or device joins the server, HAMMER performs a one-time frame alignment to bring the new robot’s data stream into alignment with the server’s map frame (Fig. 2). This computation takes approximately $36$ sec. in our implementation. During this time, the robot’s data stream is cached at the server and integrated into the map after an alignment is found. From then on, the robot’s data is continually fused into the server’s $3$DGS map in real time.
III-B1 Correspondences
@@
First, HAMMER detects potential candidate correspondences between the data of the unaligned robots and that of the aligned robots. To succinctly describe the content of images, a place-recognition feature vector is extracted for every color image received from the collaborating robots using NETVLAD [25]. The place features from the unaligned robots are periodically compared to those of the aligned robots using cosine similarity. For a pair of images derived from aligned robot $i$ and unaligned robot $j$ with similarities exceeding $\gamma_{ij}$, an additional feature-matching verification step is performed where per-image SuperPoint features [26] are extracted and matched using LightGlue [27]. If the fraction of matched features exceeds a fixed ratio $\xi=0.25$ then the image pair is accepted as a potential inter-robot correspondence.
Figure 2: HAMMER uses a one-time computation to align a new robot’s data stream with the server’s map frame. We match images from the unaligned robot with images from the server map using a place recognition feature extractor. When a correspondence is verified, matched images from robot and server are passed to an SfM solver, from which we find a single transform to align the robot to the server map frame. From then on, images and transformed SLAM poses from the robot are streamed into the data loader where they are continually refined as a part of $3$DGS training.
III-B2 Adaptive Thresholding
@@
When no image pairs exceed $\xi$, then $\gamma_{ij}$ is increased to match the mean of all place recognition similarities greater than the previous value. Adaptive thresholding helps combat unilaterally higher similarities of images captured from the same camera on different robots regardless of visual content. $\gamma_{ij}$ is initially set to a low value for all robot pairs, circumventing any hyper-parameter tuning that would be necessary. While determining candidate alignment image pairs grows quadratically in the number of images per robot, many of these operations can be cached, forming a relatively small computational footprint.
III-B3 Localized Structure-from-Motion
@@
For every potential correspondence between an aligned and unaligned robot, a small set of color images within some fixed temporal window size $W=16$ is selected from the aligned and unaligned robot’s data cache on the server (i.e. $32$ images). These images are used to perform localized SfM. This procedure is in contrast to bulk SfM, where the entirety of the robot’s data stream is used. To perform SfM, we use the COLMAP backend [18] with SuperPoint features and the SuperGlue matcher [28], which have exhibited robustness in aligning images from heterogeneous devices. Importantly, the localized SfM does not incorporate any pose information from the onboard SLAM systems, producing an independent relative pose estimate between robots.
III-B4 Registration
@@
The output of the localized SfM procedure in $\mathcal{T}^{s}$ provides a transformation link from the unaligned robot’s SLAM frame $\mathcal{T}^{j}$ to the aligned robot’s world frame $\mathcal{T}^{g,i}$, which by definition is also the global frame of all other aligned robots $\mathcal{T}^{g}$. Computing the relative transformation from $\mathcal{T}^{j}\to\mathcal{T}^{s}$ and $\mathcal{T}^{s}\to\mathcal{T}^{g}$ requires solving two instances of a modified absolute orientation problem [29] for scale, rotation, and translation. The first instance computes the transformation from unaligned robot $j$’s $W$ SLAM poses in $\mathcal{T}^{j}$ to the corresponding $W$ SfM poses in $\mathcal{T}^{s}$. The other transforms aligned robot $i$’s $W$ SfM solutions in $\mathcal{T}^{s}$ to its corresponding aligned $W$ poses in global frame $\mathcal{T}^{g}$. To compute these transformations, we solve the following rotation-aware absolute orientation problem
@@
where $||\cdot||_{F}$ denotes the Frobenius norm. Equation 1 optimizes the scaling, rotation, and translation $(s,R,t)$ between the two frames with a small regularization term on the rotation to address degenerate data. While no closed form solution exists for (1), each individual term has a closed form solution. Therefore, for small $\epsilon$, an approximately optimal solution can be computed.
The composition of both relative transformations yields $(s^{jg},R^{jg},t^{jg})$, the relative frame transformation $\mathcal{T}^{j}\to\mathcal{T}^{g}$. Importantly, given that the origin robot estimates poses in metric scaling, all $s^{jg}$ scales to a metric coordinate frame $\mathcal{T}^{g,0}$. Therefore, other than the origin robot, no other robots need to run metric SLAM methods.
The proposed inter-robot alignment module is robust and fails gracefully by terminating early or rejecting spurious alignments. During runtime, HAMMER rejects alignments where the localized SfM fails to estimate poses for all $2W$ input images or alignments that have high translation ($0.1$m in the map frame) or rotation errors ($10\deg$) as computed by (1). We also emphasize that the alignment process is performed only once for each robot with the exception of the origin robot which does not require alignment. Specifically, upon the first successful alignment, robot $j$ is immediately designated as “aligned” and no longer performs any alignment procedure even if there are available potential correspondences. Therefore, although alignment is costly, it is a one-time cost to produce the transform $\mathcal{T}^{j}\to\mathcal{T}^{g}$.
@@
Figure 2 illustrates a high-level overview of the alignment process. Importantly, the process treats the onboard localization algorithms as black-boxes and only uses the resulting camera poses and color images as input. This attribute of HAMMER dramatically simplifies the integration of new devices due to the ubiquity and variety of on-board localization frameworks. In contrast, alignment methods that require access to lower-level information from each individual robot’s localization system (e.g. pose graphs or sparse maps) will fail if the software is closed-source.
III-C Multi-robot 3DGS Mapping